For each of the following textbook problems (ch3)
Answer #1

Initial speed of golf ball in the horizontal direction is u_x=11/4\,m/s

Initial speed of golf ball  in the vertical direction is u_y=0

Height of the cliff is h=15.5\,m

a)

using S=ut+\frac{1}{2}at^2

h=u_yt+\frac{1}{2}gt^2

15.5=(0)t+\frac{1}{2}(9.8)t^2

t=1.78\,s

Time during which the ball is in air is t=1.78\,s

b)

Horizontal distance from the cliff at which the ball strikes water is   x=u_xt

x=(11.4)(1.78)=20.3\,m

c)

using v^2-u^2=2aS

v_x^2-u_x^2=0   that is v_x=u_x=11.4\,m/s

v_y^2-u_y^2=2gh   that is v_y=\sqrt{2gh}=\sqrt{2\times9.8\times15.5}=17.4\,m/s